From a thin disc mass M and radius R disc of radius R/3 is cut the centre of mass shifted from O to O_2 what is the value of o_2

Grade:11

1 Answers

jitender
6 years ago
Since mass of 4£R^2= MThen mass of 4£(R/3)^2=M/9Now using the formula..Xcm= (m1x1 + m2x2)/m1+m2O_2=(M×0 - (M/9)2R/3)/M-M/82R/3= position of centre of mass of removed disk from centre of original disk.O_2= -R/12That is centre of mass shifts R/12 towards left